$SPAD/src/input elt.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool elt.output
)set message test on
)set message auto off
)clear all

———-

This file tests elts and setelts — * —

--S 1 of 4
u : Bits := new(10,true)
--R
--R
--R   (1)  "1111111111"
--R                                                                   Type: Bits
--E 1
--S 2 of 4
u(3..5) := false; u
--R
--R
--R   (2)  "1100011111"
--R                                                                   Type: Bits
--E 2
)clear all
--S 3 of 4
u:Any := [1, 7.2, 3/2, x**2, "wally"]
--R
--R
--R               3  2
--R   (1)  [1,7.2,-,x ,"wally"]
--R               2
--R                                                              Type: List(Any)
--E 3
--S 4 of 4

u.1
--R
--R
--R   (2)  1
--R                                                        Type: PositiveInteger
--E 4
)spool
)lisp (bye)

———-

References^{[1] nothing}

5